# Wu’s Method can Boost Symbolic AI to Rival Silver Medalists and AlphaGeometry to Outperform Gold Medalists at IMO Geometry

### 09 Apr 2024 v1

@@ 1 Introduction
In this work, we put the capabilities of Wu’s method to the test on problems that involve meaningful and structured mathematical statements, rather than often arbitrary worst case instances [13]. To this end, we re-evaluate Wu’s Method on the IMO-AG-30 benchmark introduced by Trinh et al. [26]. We find that it performs surprisingly strong, solving 15 problems, and some of them are not solved by any of the other methods. This leads to two key findings:
@@ 3 Conclusions
On the other hand, the notable success of our approach suggests that IMO geometry problems, while challenging for humans, might not fully test the limits of modern computational solvers. Repeated application of human-defined heuristics combined with few constructions (up to three construction have been previously shown to be sufficient [26]) for achieving success rather than on a deep, complex exploration of combinatorial possibilities. This situation is reminiscent of chess endgames, which were mastered relatively early on by brute-force solvers [14, 21]. We hope this note encourages development of new benchmarks for automated theorem provers in geometry.
IMO Geometry feels like a chess endgame, fixed with little wiggle room.
The real fun lies in teaching machines to solve the Geometry middlegames!

### 11 Apr 2024 v2

@@ 1 Introduction
Rather than indiscriminately tackling arbitrary problem instances, mathematicians concentrate their efforts on statements exhibiting specific properties that render them interesting, meaningful, and tractable within the broader context of mathematical inquiry [13]. In this work, we put the capabilities of Wu’s method to the test on such structured problems. To this end, we re-evaluate Wu’s Method on the IMO-AG-30 benchmark introduced by Trinh et al. [24]. We find that it performs surprisingly strong, solving 15 problems, some of which are not solved by any of the other methods. This leads to two key findings:
@@ 3 Conclusions
Our exploration of the frontier between synthetic methods mimicking human reasoning processes and the capabilities of more abstract algebraic methods is motivated by the idea that the similarity to human reasoning and the generality of intelligence are distinct concepts, each with its own merits and applications. We believe that the strength of algebraic methods transcends the realm of geometry, promising significant advancements in areas as varied as compiler verification and beyond. This potential underscores our belief in the necessity to broaden the scope of challenges addressed by automated theorem proving. We envision the development of future benchmarks to strive for diversity and, potentially, open-ended testing. Embracing a wider array of problems will likely bring new insights on the usefulness, limitations, and interplay of neural and symbolic methods for general reasoning skills.
IMO Geometry may be compared with a chess endgame, fixed with little wiggle room.
We can’t wait to see machines start solving the Geometry middlegames!